WELL THE QUESTION IS (USING "A SQUARE +B SQUARE+C SQUARE" ,FIND THE LENGTH OF THE HYPOTENUSE OF THE TRIANGLE BELOW. SO, THE TRIANGLE HAS A=120 ,B=90 AND C=? SO WHAT THE ANSWER WOULD BE "C"?
1
Expert's answer
2011-10-06T10:43:50-0400
C = sqrt(A^2 + B^2) = sqrt(120^2 + 90^2) = sqsrt(14400+8100) = sqrt(22500) =
150